# Phases and phase transition in Grover’s algorithm with systematic noise

### v1

@@ IV Conclusion
We analyzed Grover’s algorithm with systematic noise by modeling the Grover operator as a Floquet unitary. We showed that the bulk energy spectrum is well described by an effective Hamiltonian whose spectral statistics match the Gaussian unitary ensemble but whose Gaussian eigenvalue distribution reflects the underlying $k$-local nature of the noise model. We use this structure to derive sharp crossovers between behaviors, finding a gap-closing ergodicity transition at relatively large noise strength $\delta_{c,\rm gap}\sim 1/L$ and a separate transition where computational power is lost above an exponentially small noise strength $\delta_{c,\rm comp}\sim L^{-1/2}2^{-L/2}$. The random matrix theory model enables analytical predictions for many quantities, such as these noise thresholds and dynamics of the initial state.
@@
The machinery we have developed for this simplest case of Grover’s algorithm with a single solution should provide insight into more generally applying many-body physics approaches to noisy quantum algorithms. For example, we can readily see that Grover’s algorithm with multiple solutions exhibits the same phenomenology, since the core aspects of a degenerate gapped bulk and cat-like special eigenstates holds there as well. A more interesting variant is to instead consider decompositions of the multiqubit Toffoli gate using ancilla qubits. If the ancillas are erased before reuse, they will behave similar to a conventional noise channel which previous results suggest will yield and exponentially small error threshold. On the other hand, if the ancillas are uncomputed and the measured and postselected, the resulting dynamics is noisy but non-unitary, similar to dynamics generated by non-Hermitian Hamiltonians. Non-Hermitian dynamics are known to have qualitatively different phase structures than their Hermitian counterparts. How this affects the phase structure and noise tolerance of Grover’s algorithm remains a fascinating open question.
@@
Another important avenue to investigate for Grover’s algorithm is the role of non-$k$-local gating. While most controllable quantum computers rely on $k$-local gates - even if swap operations or atomic repositioning relax the assumption of strict spatial locality – there are potential routes to remove this restriction by using a non-local cavity mode or the long-range tails of Rydberg blockade [35]. This not only allows more efficient decomposition of the multicontrolled gates, but also implies that systematic gate errors will be nonlocal as well. Nonlocal errors in $H_{\mathrm{eff}}$ will cause it to behave more like a proper GUE random matrix, for which a power law noise tolerance $\delta\sim 1/L^{\alpha}$ can in principle be admitted for both the ergodicity transition and the computational phase transition. In the NISQ era of finite noise, it is crucial to know these noise scalings in order to determine near-term experimental feasibility. We therefore suggest that non-local gating may be a pathway to drastically improved noise-tolerance for Grover’s algorithm, based on our analytical understanding through random matrix theory.
@@
Finally, while Grover’s algorithm provides a concrete example of Floquet analysis in the context of useful quantum algorithms, the eventual goal is to apply similar tools to a wider class of algorithms. Rather than approaching such problem on an individual basis, we would hope to find inspiration for doing so in “grand unified” theories of quantum computation [36]. While no Floquet structure is immediately apparently in such prescriptions, we might hope to find a constituent block on which similar analysis can be applied and use that insight to understand systematic noise tolerance of a wider class of quantum algorithms.
@@ V Acknowledgements
We thank Sarang Gopalakrishnan, Arjun Mirani, and Junpeng Hou for useful discussions. This work was performed with support from the National Science Foundation (NSF) through award numbers OMR-2228725 and DMR-1945529, the Welch Foundation through award number AT-2036-20200401, and DARPA through award number HR00112330022 (M.K. and S.D.). V.K. acknowledges support from the Office of Naval Research Young Investigator Program (ONR YIP) under Award Number N00014-24-1-2098, the Alfred P. Sloan Foundation through a Sloan Research Fellowship and the Packard Foundation through a Packard Fellowship in Science and Engineering. C.Z. acknowledges support from AFOSR (FA9550-20-1-0220), NSF (PHY-2409943, OMR-2228725, ECCS-2411394, OSI-2326628). Part of this work was performed at the Aspen Center for Physics, which is supported by NSF grant No. PHY-1607611, and at the Kavli Institute for Theoretical Physics, which is supported by NSF grant No. PHY-1748958.
@@ Appendix A Structure of effective Hamiltonian
We start with a heatmap of the magnitude of the elements of $H_{\rm eff}$ in the computational ($z$) basis and in the eigenbasis of $G_{0}$ in Figure 10 . A random matrix would show uniform average color in any basis, with no clear distinction between diagonal and off-diagonal matrix elements. By contrast, $H_{\mathrm{eff}}$ shows band like structures with strongest matrix elements corresponding to the diagonal. This follows from the $k$-locality of the gate perturbations in the structure of $H_{\rm eff}$ (Eq. (24)), which means that only a limited number of spins can be flipped in the $z$-basis (and the same applies to the eigenbasis of $G_{0}$).
@@ Appendix B Dynamics of the special states
In addition to gap-closing and computational phase transitions, the random matrix treatment gives analytical predictions for nearly all properties of the algorithm’s behavior. In this section, we describe the behavior of special state dynamics, which is the most relevant dynamical feature that allows the unperturbed algorithm $G_{0}$ to be successful. While we can make predictions for how systematic noise modifies these dynamics, their computational relevance is limited to the very small regime of $\delta\ll\delta_{c,\rm comp}$. However, we can also make concrete predictions for ergodicity breaking as long as $\delta<\delta_{c,\rm gap}$.
@@
$|\psi^{(0)}_{1}\rangle$ $=$ $\cos\left(\frac{\theta^{{}^{\prime}}}{2}\right)|0\rangle+e^{i\phi}\sin\left(\frac{\theta^{{}^{\prime}}}{2}\right)|\bar{x}\rangle,$ (46)
$|\psi^{(0)}_{2}\rangle$ $=$ $\sin\left(\frac{\theta^{{}^{\prime}}}{2}\right)|0\rangle-e^{i\phi}\cos\left(\frac{\theta^{{}^{\prime}}}{2}\right)|\bar{x}\rangle,$ (47)
@@
$G(\delta)|\psi_{j}\rangle=e^{-i\delta H^{\rm spec}_{\rm eff}}\tilde{G}|\psi_{j}\rangle,$ (48)
@@
While the exact expression for this off-diagonal matrix element, $|\langle\bar{x}_{k}|H^{{}^{\prime}}_{\rm eff}|\psi^{(0)}_{j}\rangle|^{2}$, depends on the particular choice of $k$, the sum over these matrix elements is constrained to be no larger than $E_{0}^{2}$ (at least upon averaging over $j$) due to the fact that $\mathrm{Tr}[H_{\mathrm{eff}}^{2}]=E_{0}^{2}N$ can be broken up into diagonal and off-diagonal pieces and the diagonals already given the $E_{0}^{2}N$ scaling. Therefore, the off-diagonal sum must give $E_{0}^{2}N$ or smaller, and the lack of summing over $j$ in Eq. 51 removes the multiplicative factor $N$. Assuming that the sum saturates this scaling, we conclude that
@@
Next, let us analyze the behavior of the probabilities $P_{0}(t)$ and $P_{\bar{x}}(t)$ in presence of noise. Consider the following ansatz for the two special states within first order in noise:
$|\psi_{1}\rangle$ $=$ $\left(1-P_{b}\right)^{1/4}\left[\cos\left(\frac{\theta^{\prime}}{2}\right)|\bar{x}\rangle+e^{i\phi^{\prime}}\sin\left(\frac{\theta^{\prime}}{2}\right)|0\rangle\right]$
@@
$|\psi_{2}\rangle$ $=$ $\left(1-P_{b}\right)^{1/4}\left[\sin\left(\frac{\theta^{\prime}}{2}\right)|\bar{x}\rangle-e^{i\phi^{\prime}}\cos\left(\frac{\theta^{\prime}}{2}\right)|0\rangle\right]$
@@
where we will see that $P_{b}$ is a related parameter that describes late time occupation of the bulk given our actual starting state. This is an unusual ansatz, as it dresses the zeroth order result from degenerate perturbation theory by a non-perturbative overall rescaling of the amplitude, $\left(1-P_{b}\right)^{1/4}$. This comes from the finite “leakage” into the bulk at late times but, as we’ll show, seems to give a self-consistent picture of the full dynamics.
@@
Having decomposed the initial state into eigenstates (Eq. 53), we can now easily calculate its time evolution as
@@
Term A: Late-time average: The existence of a non-exponentially-small occupation in state $|\bar{x}\rangle$ is a clear indication of nonergodicity. It will differ between noise realizations, but we can estimate its noise average (which is equal to participation ratio in the eigenstate basis):
@@
$\approx\left(1-P_{b}\right)\left[\cos^{4}\left(\frac{\theta^{\prime}}{2}\right)+\sin^{4}\left(\frac{\theta^{\prime}}{2}\right)\right]+O(1/N)$
@@
Term C: Damped oscillations from bulk continuum: Due to the $\pi$ gap between bulk and special states, there are fast (period 2) oscillations in the occupation. However, since the bulk has finite spectral width, these oscillations are damped. The gap $\Delta$ between $E_{1}$ and $E_{2}$ will not be resolvable on the decay time-scale, so for all intents and purposes we can just set both of these to $\pi$. Using the Gaussian density of states obtained in the main text,
@@
which is simply Gaussian relaxation on time scale $\tau\sim 1/\left(\delta E_{0}\right)\sim\delta_{c,\rm gap}/\delta$ that we expect to see. Notice that this is a fast $O(1)$ time scale as long we keep $\delta/\delta_{c,\rm gap}$ fixed as we take $L$ large.
Term D: Bulk decay: Finally, the last term tells us how Loschmidt echo of a given state – in this case $|\bar{x}\rangle$ – decays when propagated by the random bulk matrix $H_{\rm eff}$. Again, using the Gaussian approximation for density of states, we find
@@
This is similar to the decay rate of the $2T$-periodic oscillations (term C) with a rate that is $\sqrt{2}$ times faster. Lastly, as a sanity check, let us confirm that $P_{\bar{x}}(0)=1$:
@@
The predicted behaviors above are seen in Figure 11. Note that the period-$2$ oscillations do appear on close inspection of the short-time behavior (cf. the sharp jump in the curve for $\delta=0.17$ in orange), but are not clear on the longer time scale plotted here.
Figure 11: Plot of average probability $P_{0}(t)$ and $P_{\bar{x}}(t)$ for $L=10$ as a function of time steps for different noise strengths: $0.170.01<\delta_{c,\rm comp}\approx 0.02<0.06<0.08<0.1<\delta_{c,\rm gap}\approx 0.15<$ 0.17. When $\delta<\delta_{c,\rm gap}$, the probabilities oscillates with a larger frequency but smaller amplitude as compared to the noiseless case. The noises $\delta=0.06,0.08,0.1$ lie in between $\delta_{c,\rm gap}$ and $\delta_{c,\rm comp}$, for which there is no significant change in the probabilities as anticipated from the above analytical calculation. Noise $\delta=0.17>\delta_{c,\rm comp}$ represents the noise dominated dynamics. Error bands represents the standard error of mean for $256$ noise realizations.

### v2

@@ IV Conclusion
We analyzed Grover’s algorithm with systematic noise by modeling the Grover operator as a Floquet unitary. We showed that the bulk energy spectrum is well-described by an effective Hamiltonian whose spectral statistics match the Gaussian unitary ensemble but whose Gaussian eigenvalue distribution reflects the underlying $k$-local nature of the noise model. We use this structure to derive sharp crossovers between behaviors, finding a gap-closing ergodicity transition at relatively large noise strength $\delta_{c,\rm gap}=\mathcal{O}(L^{-1})$ and a separate transition where computational power is lost above an exponentially small noise strength $\delta_{c,\rm comp}=\mathcal{O}(2^{-L/2})$. The random matrix theory model enables analytical predictions for many quantities, such as these noise thresholds and dynamics of the initial state.
@@
The machinery we have developed for this simplest case of Grover’s algorithm with a single solution should provide insights into more generally applying many-body physics approaches to noisy quantum algorithms. For example, we can readily see that Grover’s algorithm with multiple solutions exhibits the same phenomenology since the core aspects of a degenerate gapped bulk and cat-like special eigenstates hold there as well. While we specifically focus on unraveling both the oracle and Grover diffusion operator, the same effect will occur if we just unravel the Grover’s diffusion operator (reflection about $|x\rangle$) since it is identical to the oracle for a single target state. Introducing noise to only one of these operators, rather than both, would merely adjust the critical disorder thresholds by approximately a factor of 2, affecting only the scaling prefactor rather than the fundamental behavior. A more interesting variant is to instead consider decompositions of the multiqubit Toffoli gate using ancilla qubits. If the ancillas are erased before reuse, they will behave similar to a conventional noise channel which previous results suggest will yield an exponentially small error threshold. On the other hand, if the ancillas are uncomputed and the measured and postselected, the resulting dynamics is noisy but nonunitary, similar to dynamics generated by non-Hermitian Hamiltonians. Non-Hermitian dynamics are known to have qualitatively different phase structures than their Hermitian counterparts. How this affects the phase structure and noise tolerance of Grover’s algorithm remains a fascinating open question.
@@
Another important avenue to investigate for Grover’s algorithm is the role of non-$k$-local gating. While most controllable quantum computers rely on $k$-local gates - even if swap operations or atomic repositioning relax the assumption of strict spatial locality – there are potential routes to remove this restriction by using a nonlocal cavity mode or the long-range tails of Rydberg blockade [38]. This not only allows more efficient decomposition of the multicontrolled gates, but also implies that systematic gate errors will be nonlocal as well. Nonlocal errors in $H_{\mathrm{eff}}$ will cause it to behave more like a proper GUE random matrix, for which a power law noise tolerance $\delta\sim 1/L^{\alpha}$ can in principle be admitted for both the ergodicity transition and the computational transition. In the NISQ era of finite noise, it is crucial to know these noise scalings in order to determine near-term experimental feasibility. We therefore suggest that nonlocal gating may be a pathway to drastically improved noise tolerance for Grover’s algorithm, based on our analytical understanding through random matrix theory.
In the context of many-body Floquet systems, decorrelating unitary errors between Floquet cycles often has the effect of destabilizing the phase of matter rather than slowing its decay. For example, in the topological Floquet-Anderson insulator, one of us showed in previous work [39] that it is stable to weak systematic noise, whereas uncorrelated noise between Floquet cycles leads to eventual decay. A similar situation occurs in many other Floquet phases of matter, such as time crystals, which are absolutely stable to weak systematic perturbations as one of us showed in [40] but unstable to nonsystematic noise [41]. However, our analysis suggest that the application of twirling operations to decorrelate the noise could significantly mitigate the impact of errors, improving the scaling from $2^{-L/2}$ to $2^{-L/4}$. This observation is consistent with the results of Shapira et al. [12], although their analysis did not fully consider the effect of unitary noise in the oracle or the Grover diffusion operator.
@@
Finally, while Grover’s algorithm provides a concrete example of Floquet analysis in the context of useful quantum algorithms, the eventual goal is to apply similar tools to a wider class of algorithms. Rather than approaching such problem on an individual basis, we would hope to find inspiration for doing so in “grand unified” theories of quantum computation [42]. While no Floquet structure is immediately apparent in such prescriptions, we might hope to find a constituent block on which similar analysis can be applied and use that insight to understand systematic noise tolerance of a wider class of quantum algorithms.
@@ V Acknowledgements
We thank Sarang Gopalakrishnan, Arjun Mirani, and Junpeng Hou for useful discussions. This work was performed with support from the National Science Foundation (NSF) through Awards No. OMR-2228725 and DMR-1945529, the Welch Foundation through Award No. AT-2036-20200401, and DARPA through Award No. HR00112330022 (M.K. and S.D.). V.K. acknowledges support from the Office of Naval Research Young Investigator Program (ONR YIP) under Award No. N00014-24-1-2098, the Alfred P. Sloan Foundation through a Sloan Research Fellowship and the Packard Foundation through a Packard Fellowship in Science and Engineering. C.Z. acknowledges support from AFOSR (FA9550-20-1-0220), NSF (PHY-2409943, OMR-2228725, ECCS-2411394, OSI-2326628). Part of this work was performed at the Aspen Center for Physics, which is supported by NSF Grant No. PHY-1607611, and at the Kavli Institute for Theoretical Physics, which is supported by NSF Grant No. PHY-1748958.
@@ Appendix A Structure of effective Hamiltonian
We start with a heatmap of the magnitude of the elements of $H_{\rm eff}$ in the computational ($z$) basis and in the eigenbasis of $G_{0}$ in Fig 10. A random matrix would show uniform average color in any basis, with no clear distinction between diagonal and off-diagonal matrix elements. By contrast, $H_{\mathrm{eff}}$ shows band-like structures with the strongest matrix elements corresponding to the diagonal. This follows from the $k$-locality of the gate perturbations in the structure of $H_{\rm eff}$ (Eq. (24)), which means that only a limited number of spins can be flipped in the $z$-basis (and the same applies to the eigenbasis of $G_{0}$).
@@ Appendix B Dynamics of the special states
In addition to gap-closing and computational transitions, the random matrix treatment gives analytical predictions for nearly all properties of the algorithm’s behavior. In this appendix, we describe the behavior of special state dynamics, which is the most relevant dynamical feature that allows the unperturbed algorithm $G_{0}$ to be successful. While we can make predictions for how systematic noise modifies these dynamics, their computational relevance is limited to the very small regime of $\delta\ll\delta_{c,\rm comp}$. However, we can also make concrete predictions for ergodicity breaking as long as $\delta<\delta_{c,\rm gap}$.
@@
$|\psi^{(0)}_{1}\rangle$ $=$ $\cos\left(\frac{\theta^{{}^{\prime}}}{2}\right)|\textbf{0}\rangle+e^{i\phi}\sin\left(\frac{\theta^{{}^{\prime}}}{2}\right)|\bar{x}\rangle,$ (46)
$|\psi^{(0)}_{2}\rangle$ $=$ $\sin\left(\frac{\theta^{{}^{\prime}}}{2}\right)|\textbf{0}\rangle-e^{i\phi}\cos\left(\frac{\theta^{{}^{\prime}}}{2}\right)|\bar{x}\rangle,$ (47)
@@
$G(\delta)|\psi_{j}\rangle=e^{-i\delta H^{\rm spec}_{\rm eff}}\tilde{G}|\psi_{j}\rangle;$ (48)
@@
While the exact expression for this off-diagonal matrix element, $|\langle\bar{x}_{k}|H^{{}^{\prime}}_{\rm eff}|\psi^{(0)}_{j}\rangle|^{2}$, depends on the particular choice of $k$, the sum over these matrix elements is constrained to be no larger than $E_{0}^{2}$ (at least upon averaging over $j$) due to the fact that $\mathrm{Tr}[H_{\mathrm{eff}}^{2}]=E_{0}^{2}N$ can be broken up into diagonal and off-diagonal pieces and the diagonals already given the $E_{0}^{2}N$ scaling. Therefore, the off-diagonal sum must give $E_{0}^{2}N$ or smaller, and the lack of summing over $j$ in Eq.( 51) removes the multiplicative factor $N$. Assuming that the sum saturates this scaling, we conclude that
@@
Next, let us analyze the behavior of the probabilities $P_{0}(t)$ and $P_{\bar{x}}(t)$ in the presence of noise. Consider the following ansatz for the two special states within first order in noise:
$|\psi_{1}\rangle$ $=$ $\left(1-P_{b}\right)^{1/4}\left[\cos\left(\frac{\theta^{\prime}}{2}\right)|\bar{x}\rangle+e^{i\phi^{\prime}}\sin\left(\frac{\theta^{\prime}}{2}\right)|\textbf{0}\rangle\right]$
@@
$|\psi_{2}\rangle$ $=$ $\left(1-P_{b}\right)^{1/4}\left[\sin\left(\frac{\theta^{\prime}}{2}\right)|\bar{x}\rangle-e^{i\phi^{\prime}}\cos\left(\frac{\theta^{\prime}}{2}\right)|\textbf{0}\rangle\right]$
@@
where we will see that $P_{b}$ is a related parameter that describes late time occupation of the bulk given our actual starting state. This is an unusual ansatz, as it dresses the zeroth order result from degenerate perturbation theory by a nonperturbative overall rescaling of the amplitude, $\left(1-P_{b}\right)^{1/4}$. This comes from the finite “leakage” into the bulk at late times but, as we’ll show, seems to give a self-consistent picture of the full dynamics.
@@
Having decomposed the initial state into eigenstates [Eq.( 53)], we can now easily calculate its time evolution as
@@
Term A: Late-time average: The existence of a non-exponentially-small occupation in state $|\bar{x}\rangle$ is a clear indication of nonergodicity. It will differ between noise realizations, but we can estimate its noise average (which is equal to the participation ratio in the eigenstate basis):
@@
$\approx\left(1-P_{b}\right)\left[\cos^{4}\left(\frac{\theta^{\prime}}{2}\right)+\sin^{4}\left(\frac{\theta^{\prime}}{2}\right)\right]+O(1/N).$
@@
Term C: Damped oscillations from bulk continuum: Due to the $\pi$ gap between bulk and special states, there are fast (period 2) oscillations in the occupation. However, since the bulk has finite spectral width, these oscillations are damped. The gap $\Delta$ between $E_{1}$ and $E_{2}$ will not be resolvable on the decay time scale, so for all intents and purposes we can just set both of these to $\pi$. Using the Gaussian density of states obtained in the main text,
@@
which is simply Gaussian relaxation on the time scale $\tau\sim 1/\left(\delta E_{0}\right)\sim\delta_{c,\rm gap}/\delta$ that we expect to see. Notice that this is a fast $O(1)$ time scale as long we keep $\delta/\delta_{c,\rm gap}$ fixed as we take $L$ large.
Term D: Bulk decay: Finally, the last term tells us how Loschmidt echo of a given state – in this case $|\bar{x}\rangle$ – decays when propagated by the random bulk matrix $H_{\rm eff}$. Again, using the Gaussian approximation for the density of states, we find
@@
This is similar to the decay rate of the $2T$-periodic oscillations (term C) with a rate that is $\sqrt{2}$ times faster. Finally, as a sanity check, let us confirm that $P_{\bar{x}}(0)=1$:
@@
The predicted behaviors above are seen in Fig 11. Note that the period-$2$ oscillations do appear on close inspection of the short-time behavior (cf. the sharp jump in the curve for $\delta=0.17$ in orange), but are not clear on the longer time scale plotted here.
Figure 11: Plot of average probability $P_{0}(t)$ and $P_{\bar{x}}(t)$ for $L=10$ as a function of time steps for different noise strengths: $0.170.01<\delta_{c,\rm comp}\approx 0.02<0.06<0.08<0.1<\delta_{c,\rm gap}\approx 0.15<$ 0.17. When $\delta<\delta_{c,\rm gap}$, the probabilities oscillate with a larger frequency but smaller amplitude as compared to the noiseless case. The noises $\delta=0.06,0.08,0.1$ lie between $\delta_{c,\rm gap}$ and $\delta_{c,\rm comp}$, for which there is no significant change in the probabilities as anticipated from the above analytical calculation. Noise $\delta=0.17>\delta_{c,\rm comp}$ represents the noise dominated dynamics. Error bands represents the standard error of mean for $256$ noise realizations.